OpenStudy (anonymous):

integral of (sin^2 (x))

OpenStudy (anonymous):

integral, not derivative please

OpenStudy (anonymous):

x/2-1/4 sin(2 x)+c

OpenStudy (anonymous):

Write sin^2(x) as 1/2-1/2 cos(2 x): = integral (1/2-1/2 cos(2 x)) dx Integrate the sum term by term and factor out constants: = integral 1/2 dx-1/2 integral cos(2 x) dx For the integrand cos(2 x), substitute u = 2 x and du = 2 dx: = integral 1/2 dx-1/4 integral cos(u) du The integral of cos(u) is sin(u): = integral 1/2 dx-(sin(u))/4 The integral of 1/2 is x/2: = x/2-(sin(u))/4+constant Substitute back for u = 2 x: = x/2-1/4 sin(2 x)+constant Which is equal to: = 1/2 (x-sin(x) cos(x))+constant